# Beyond Reward Hacking: Causal Rewards for Large Language Model Alignment

### v1

@@ Abstract
Recent advances in large language models (LLMs) have demonstrated significant progress in performing complex tasks. While Reinforcement Learning from Human Feedback (RLHF) has been effective in aligning LLMs with human preferences, it is susceptible to spurious correlations in reward modeling. Consequently, it often introduces biases—such as length bias, sycophancy, conceptual bias, and discrimination—that hinder the model’s ability to capture true causal relationships. To address this, we propose a novel causal reward modeling approach that integrates causal inference to mitigate these spurious correlations. Our method enforces counterfactual invariance, ensuring reward predictions remain consistent when irrelevant variables are altered. Through experiments on both synthetic and real-world datasets, we show that our approach mitigates various types of spurious correlations effectively, resulting in more reliable and fair alignment of LLMs with human preferences. As a drop-in enhancement to the existing RLHF workflow, our causal reward modeling provides a practical way to improve the trustworthiness and fairness of LLM finetuning.
@@ 1 Introduction
To understand and mitigate these issues, it is essential to consider the sources of error in reward modeling. The total error in the reward model can be decomposed into reducible and irreducible components, as shown in Eq. (1).
Total $\text{Total Error}=\underbrace{\text{Bias}^{2}+\text{Variance}}_{\text{Reducible Error}}+\underbrace{\text{Noise}}_{\text{Irreducible Error}}$ (1)
The reducible error comprises estimation errors stemming from limited data and model approximation, which can be alleviated by collecting more data or increasing model capacity. However, irreducible error originates from inherent noise and imperfections in the data, such as the spurious correlations described earlier, which cannot be resolved merely by increasing data quantity or model complexity (Geman et al., 1992). For instance, if longer responses are disproportionately represented and favored among higher-reward examples, the reward model may learn to prefer longer outputs irrespective of their quality, leading to the length bias observed in RLHF policies. Similarly, human annotators may unintentionally favor responses that flatter them. This bias can mislead the model, causing it to prefer agreeableness over truthfulness (Perez et al., 2022). Notably, such biases cannot be mitigated by simply increasing the size of the dataset. On the contrary, it may further exacerbate the effects of reward hacking (Ribeiro et al., 2016).
To address this challenge, we propose a novel approach in this work that integrates causality into reward modeling to mitigate the impact of spurious correlations and prevent reward hacking in RLHF. By leveraging causal inference techniques, we develop a causal reward model (CRM) that is robust to these spurious correlations and captures the true causal relationship of responses on human preferences. Central to our method is the concept of counterfactual invariance, which ensures that the reward model’s predictions remain consistent under interventions on irrelevant aspects of the input, thereby reducing the irreducible error caused by spurious correlations (Veitch et al., 2021).
@@
  • We introduce a causal framework for reward modeling that incorporates causal regularization into the training process, allowing the model to learn “true” [^1] Here, “true” represents the user’s belief about what is true. causality from spurious relationship.
@@
  • Our method is simple to implement and can be seamlessly integrated into existing RLHF pipelines, providing a practical solution to enhance the reliability of LLMs.
@@ 2.1 Reward Hacking and Spurious Correlation
LLM reward hacking often stems from the model’s reliance on spurious correlations in the preference dataset, such as length (Sountsov and Sarawagi, 2016; Dubois et al., 2024a; Huang et al., 2024), sycophancy (Sharma et al., 2023; Ranaldi and Pucci, 2023), conceptual (Zhou et al., 2023), and demographic (Salinas et al., 2023) biases. These spurious correlations, closely linked to reward hacking, can impair a model’s capability to learn and generalize to broader scenarios (Ribeiro et al., 2016; Geirhos et al., 2020; Chen et al., 2024d).
Without proper constraints, models will exploit all available informative features during training, including unreliable spurious ones, which results in reward hacking, even if the task is very simple (Nagarajan et al., 2020; McMilin, 2022; Chen et al., 2024e). To address this, our approach integrates causal regularization into reward modeling, enabling LLMs to learn true causal relationships, mitigate the effects from spurious correlations and thereby prevent reward hacking.
@@ 2.2 Alleviating Spurious Correlations
Early efforts to mitigate spurious correlations and reward hacking in RLHF have primarily focused on penalizing specific biases within reward models (Mnih et al., 2015), especially correcting for length bias. For example, Singhal et al. (2023) reveals that length-based biases in reward models significantly influence RLHF outcomes, often overshadowing non-length-related features, and proposes mitigation strategies such as balanced preference datasets, reward data augmentation, confidence-based truncation, increased KL penalties, explicit length penalties, omitting long outputs, and focusing on non-length reward metrics.
To address the overemphasis on longer response, Shen et al. (2023) proposed a Product-of-Experts (PoE) framework to decouple reward modeling from sequence length, thereby reducing the reward model’s preference for verbose but low-quality responses. Building on this, Eisenstein et al. (2023) introduced reward model ensembles to moderate reward hacking by diversifying the sources of feedback and reducing reliance on any single reward model’s spurious correlations. However, this method only partially mitigates the problem and falls short of fully eliminating reward hacking. More recently, Ramé et al. (2024) proposed Weight Averaged Reward Models (WARM), which enhance robustness to distribution shifts by averaging model weights, offering a more efficient and effective alternative to ensemble-based policy interpolation. ODIN (Chen et al., 2024a) advanced this line of work by introducing a disentangled reward model architecture to tackle length bias. Their approach separates reward factors into two linear heads, isolating content quality for use during RL fine-tuning, thus improving performance without sacrificing efficiency.
@@ 3.1 Reinforcement Learning from Human Feedbacks (RLHF)
Supervised finetuning (SFT).
The supervised fine-tuning step typically starts with a pre-trained language model, which is then fine-tuned through supervised learning on a high-quality dataset tailored to specific downstream tasks, such as dialogue (Bai et al., 2022), instruction following (Longpre et al., 2023), and summarization (Zheng et al., 2024). This fine-tuning process produces a model denoted as $\pi^{\text{SFT}}$.
Reward model learning.
During this stage, we will first need to have a dataset that consists of preference pairs of responses, $(y_{1},y_{2})$, for each prompt $x$. Typically, these pairs are obtained by presenting them to labelers (e.g., humans), who evaluate the responses based on their preferences, represented as $y_{w}\succ y_{l}\mid x$, where $y_{w}$ and $y_{l}$ denote the preferred and less preferred responses, respectively. From a modeling perspective, these preferences are assumed to be generated from an unknown latent reward model, $r^{*}(y,x)$. In practice, the modeling assumptions for the preferences can vary depending on the problem, but the Bradley-Terry (BT) model is a commonly used assumption. The BT model computes the probability of one response $y_{1}$ being preferred over the other response $y_{2}$ under the true reward function $r^{\star}(x,y)$ by:
@@
Fine-tuning with reinforcement learning.
Once the reward model, which serves as a proxy for the utility we aim to maximize, is trained, the next step is to apply reinforcement learning under this reward model. Typically, the following objective is used:
@@ 3.2 Counterfactual Invariance
An ideal debiased reward model should intuitively remain invariant to spurious factors of variations. For example, to eliminate length bias, the reward model should exhibit invariance to changes in response length. To formalize this notion, we leverage the concept of counterfactual invariance (Veitch et al., 2021).
We begin by introducing some notation. Let $Z$ represent the random variable corresponding to a spurious factor of variation (e.g., length), and let $T$ denote the random variable that encompasses the prompt-response pair. A reward model $r$ is said to exhibit counterfactual invariance to $Z$ if $r(T(z))=r(T(z^{\prime}))$ for all $z,z^{\prime}$, where $z$ and $z^{\prime}$ are realizations of $Z$, and $T(z)$ denotes the counterfactual $T$ we would have observed if $Z$ were $z$. Throughout the paper, we use the term “invariant" or “debiased" to refer specifically to counterfactual invariance.
@@ 3.3 Causal Decomposition
The prompt-response pair $T$ can be decomposed into latent components based on their relations with the spurious factor $Z$ (Veitch et al., 2021). Specifically, we define $T^{Z,\perp}$ as the component of $T$ that is not causally influenced by $Z$. In other words, $T^{Z,\perp}$ represents the part of $T$ such that any function of $T$ is counterfactually invariant to $Z$ if and only if it depends solely on $T^{Z,\perp}$.
Under weak conditions on $Z$, $T^{Z,\perp}$ is well defined. Further details regarding the derivation and properties can be found in (Veitch et al., 2021). In the next section, we extend this concept to develop a reward model that incorporates counterfactual invariance, which enables debiasing against various spurious factors.
@@ 4 Method
This independence condition is merely a necessary condition implied by counterfactual invariance. But the key idea is that it constrains the hypothesis class, potentially guiding the model toward learning an invariant predictor.
[…]
@@ 4.1 Maximum Mean Discrepancy (MMD) Regularization for Independence
where $\mathcal{F}$ denotes a class of functions in $\mathcal{H}_{k}$, and $x\sim\mathbb{P}$, $y\sim\mathbb{Q}$ are two random variables. Intuitively, MMD measures the maximum mean difference between $\mathbb{P}$ and $\mathbb{Q}$ over functions $f\in\mathcal{F}$, as determined by a kernel $k(\cdot,\cdot)$ such as Gaussian kernels.
In our approach, we use MMD as a regularizer to ensure that the learned reward model $f(T)$ is invariant to the spurious variable $Z$. If $Z$ is binary, our MMD regularizer can be defined as:
@@
To regularize $r_{\phi}(x,y)$, we map all responses into $M$ bins based on their spurious factor $Z$ (e.g., reponse length). For each bin $b$, we compute the conditional distribution $P_{b}(f(x,y))$. The overall objective function, combining the reward model training loss and the MMD-based regularizer, is:
@@ 5.1 Addressing Sycophantic Bias (Semi-synthetic)
Dataset and training.
To investigate sycophantic bias, we create a semi-synthetic dataset based on dataset developed by Sharma et al. (2023). Specifically, our prompts are structured with the template "{question} I think the answer is {correct_answer} but I’m really not sure.”.
In this setup, we artificially induce a correlation between sycophantic behavior and correctness. Specifically, with an $80\%$ probability, the chosen response is prefixed with "Yes, you are right." Conversely, with a $20\%$ probability, this prefix appears in the rejected response. This creates an artificial but controlled spurious correlation between agreement ("Yes, you are right.") and correct answer, enabling us to observe, measure and address sycophantic bias effectively.
@@
Results.
For each test prompt, we generate 50 responses. We then quantify sycophancy by checking whether the phrase "Yes, you are right." appears in any of those responses. Table 1 reports the percentage of test prompts for which all 50 sampled responses exhibit sycophantic behavior. It is worth noting that the SFT model, trained on chosen responses with high correlation with sycophantic phrasing, naturally tends to produce "Yes, you are right." as a default pattern. In contrast, both the conditional and unconditional CRM approaches successfully disentangle this spurious correlation and reduce the prevalence of sycophantic responses.
@@ 5.2 Addressing Length Bias
Dataset and training.
We adopted the Alpaca dataset (Dubois et al., 2024b) for our experiments. Initially, we uses the chosen response for each prompt to do supervised finetuning (SFT) using the Llama-3 8B base model (Dubey et al., 2024). Then, this SFT model was subsequently employed to train both the reward model and the policy model. For reward model, we used the chosen and the rejected pair for training. With the reward model, we then trained the SFT policy with the PPO implementation from OpenRLHF (Hu et al., 2024) for one epoch. Additional details on hyperparameters and configurations are available in the Appendix 8.2.
Results.
Our findings are illustrated in Fig. 2, where each dot on the plots represents a single model run, evaluated by its win rate, calculated as the proportion of wins against the SFT model. The score is defined by $\text{score}=50+(n_{\text{win}}-n_{\text{lose}})/N*100$, where $n_{\text{win}}$ and $n_{\text{lose}}$ denote the counts of wins and losses, respectively, and $N$ represents the total test count. In the leftmost plot, we observe that both the conditional and unconditional causal regularization methods achieve superior performance compared to the vanilla reward model with length penalty, as shown by their higher exponential moving average (EMA) curves. Furthermore, when examining the Pareto frontier, our approach demonstrates an advantage over the baseline method.
@@ Results.
Figure 2: Results on Length Bias, where each dot represents models trained with different regularization coefficients and PPO hyperparameters. The leftmost figure displays the results as an exponential moving average (EMA) curve, the middle plot illustrates the Pareto front, and the rightmost figure shows the correlation between length and rank based on reward values for different causal reward models.
@@ 5.3 Addressing Concept Bias
Dataset and training.
We conducted experiments using Yelp (Zhang et al., 2015), IMDB (Maas et al., 2011), and Amazon Shoe Review (He and McAuley, 2016) datasets, augmented with additional concept labels provided by Zhou et al. (2023). Specifically, each dataset includes three concepts, where Yelp has "price", "service", "food"; IMDB has "music", "acting", "comedy"; and Amazon has "size", "color" and "style". To introduce more obvious concept bias, following Zhou et al. (2023), we modified each dataset to ensure all positive-sentiment samples were explicitly linked to a specific concept. For instance, in the Yelp dataset, we filtered reviews so that all positive sentiment entries were linked to the "food" concept.
@@
Metrics.
We assess performance using both utility metrics (Acc@C, Acc@NoC) as well as the bias-specific metric Bias@C, as introduced in (Zhou et al., 2023). The utility metrics, which reflect the accuracy of correct sentiment classifications with (Acc@C) and without (Acc@NoC) the presence of a concept, indicate better performance with higher values. On the other hand, Bias@C measures spurious correlations associated with concept $C$, where values closer to zero suggest weaker biases. Specifically, positive Bias@C values suggest the model tends to predict positive labels when concept $C$ is present in the input, whereas negative values suggest the opposite tendency. For a more detailed explanation of the Bias@C metric, we direct interested readers to their original work (Zhou et al., 2023).
@@
Results.
As shown in Table 2, the results demonstrate that CRM consistently reduces concept bias across the Yelp, IMDB, and Amazon Shoe Review datasets compared to the vanilla reward model. Specifically, both conditional and unconditional CRMs achieve significantly lower Bias@C values, with conditional CRM showing reductions of up to 97% on the Yelp dataset (e.g., for the "Price" concept). These results highlight the effectiveness of our approach in mitigating spurious correlations.

### v2

@@ Abstract
Recent advances in large language models (LLMs) have demonstrated significant progress in performing complex tasks. While Reinforcement Learning from Human Feedback (RLHF) has been effective in aligning LLMs with human preferences, it is susceptible to spurious correlations in reward modeling. Consequently, it often introduces biases—such as length bias, sycophancy, conceptual bias, and discrimination—that hinder the model’s ability to capture true causal relationships. To address this, we propose a novel causal reward modeling approach that integrates causality to mitigate these spurious correlations. Our method enforces counterfactual invariance, ensuring reward predictions remain consistent when irrelevant variables are altered. Through experiments on both synthetic and real-world datasets, we show that our approach mitigates various types of spurious correlations effectively, resulting in more reliable and fair alignment of LLMs with human preferences. As a drop-in enhancement to the existing RLHF workflow, our causal reward modeling provides a practical way to improve the trustworthiness and fairness of LLM finetuning.
@@ 1 Introduction
To understand and mitigate these issues, it is essential to consider the sources of error in reward modeling. The total error in the reward model can be decomposed into reducible and irreducible components. The reducible error comprises estimation errors stemming from limited data and model approximation, which can be alleviated by collecting more data or increasing model capacity. However, irreducible error originates from inherent noise and imperfections in the data, such as the spurious correlations described earlier, which cannot be resolved merely by increasing data quantity or model complexity (Geman et al., 1992). For instance, if longer responses are disproportionately represented and favored among higher-reward examples, the reward model may learn to prefer longer outputs irrespective of their quality, leading to the length bias observed in RLHF policies. Similarly, human annotators may unintentionally favor responses that flatter them. This bias can mislead the model, causing it to prefer agreeableness over truthfulness (Perez et al., 2022). Notably, such biases cannot be mitigated by simply increasing the size of the dataset. On the contrary, it may further exacerbate the effects of reward hacking (Ribeiro et al., 2016).
To address this challenge, we propose a novel approach in this work that integrates causality into reward modeling to mitigate the impact of spurious correlations and prevent reward hacking in RLHF. By leveraging techniques from causality, we develop a causal reward model (CRM) that is robust to these spurious correlations and captures the true causal relationship of responses on human preferences. Central to our method is the concept of counterfactual invariance, which ensures that the reward model’s predictions remain consistent under interventions on irrelevant aspects of the input, thereby reducing the irreducible error caused by spurious correlations (Veitch et al., 2021).
@@
  • We introduce a causal framework for reward modeling that incorporates causal regularization into the training process, allowing the model to learn true [^1] Here, true represents the user’s belief about what is true. causality from spurious relationship.
@@
  • CRM is simple to implement and can be seamlessly integrated into existing RLHF pipelines, providing a practical solution to enhance the reliability of LLMs.
@@ 2.1 Reward Hacking and Spurious Correlation
LLM reward hacking often stems from the model’s reliance on spurious correlations in the preference dataset, such as length (Sountsov and Sarawagi, 2016; Dubois et al., 2024a; Huang et al., 2024), sycophancy (Sharma et al., 2023; Ranaldi and Pucci, 2023), conceptual (Zhou et al., 2023), and demographic (Salinas et al., 2023) biases. These spurious correlations, closely linked to reward hacking, can impair a model’s capability to learn and generalize to broader scenarios (Ribeiro et al., 2016; Geirhos et al., 2020; Chen et al., 2024d). Without proper constraints, models will exploit all available informative features during training, including unreliable spurious ones, which results in reward hacking, even if the task is very simple (Nagarajan et al., 2020; McMilin, 2022; Chen et al., 2024e). To address this, our approach integrates causal regularization into reward modeling, enabling LLMs to learn true causal relationships, mitigate the effects from spurious correlations and thereby prevent reward hacking.
@@ 2.2 Alleviating Spurious Correlations
Early efforts to mitigate spurious correlations and reward hacking in RLHF have primarily focused on penalizing specific biases within reward models (Mnih et al., 2015), especially correcting for length bias. For example, Singhal et al. (2023) reveals that length-based biases in reward models significantly influence RLHF outcomes, often overshadowing non-length-related features, and proposes mitigation strategies such as balanced preference datasets, reward data augmentation, confidence-based truncation, increased KL penalties, explicit length penalties, omitting long outputs, and focusing on non-length reward metrics. To address the overemphasis on longer response, Shen et al. (2023) proposed a Product-of-Experts (PoE) framework to decouple reward modeling from sequence length, thereby reducing the reward model’s preference for verbose but low-quality responses. Building on this, Eisenstein et al. (2023) introduced reward model ensembles to moderate reward hacking by diversifying the sources of feedback and reducing reliance on any single reward model’s spurious correlations. However, this method only partially mitigates the problem and falls short of fully eliminating reward hacking. More recently, Ramé et al. (2024) proposed Weight Averaged Reward Models (WARM), which enhance robustness to distribution shifts by averaging model weights, offering a more efficient and effective alternative to ensemble-based policy interpolation. ODIN (Chen et al., 2024a) advanced this line of work by introducing a disentangled reward model architecture to tackle length bias. Their approach separates reward factors into two linear heads, isolating content quality for use during RL fine-tuning, thus improving performance without sacrificing efficiency.
@@ 3.1 Reinforcement Learning from Human Feedbacks (RLHF)
Supervised fine-tuning (SFT). The SFT step typically starts with a pre-trained language model, which is then fine-tuned through supervised learning on a high-quality dataset tailored to specific downstream tasks, such as dialogue (Bai et al., 2022), instruction following (Longpre et al., 2023), and summarization (Zheng et al., 2024). This fine-tuning process produces a model denoted as $\pi^{\text{SFT}}$.
Reward model learning. During this stage, we will first need to have a dataset that consists of preference pairs of responses, $(y_{1},y_{2})$, for each prompt $x$. Typically, these pairs are obtained by presenting them to labelers (e.g., humans), who evaluate the responses based on their preferences, represented as $y_{w}\succ y_{l}\mid x$, where $y_{w}$ and $y_{l}$ denote the preferred and less preferred responses, respectively. From a modeling perspective, these preferences are assumed to be generated from an unknown latent reward model, $r^{*}(y,x)$. In practice, the modeling assumptions for the preferences can vary depending on the problem, but the Bradley-Terry (BT) model is a commonly used assumption. The BT model computes the probability of one response $y_{1}$ being preferred over the other response $y_{2}$ under the true reward function $r^{\star}(x,y)$ by:
@@
Fine-tuning with reinforcement learning. Once the reward model, which serves as a proxy for the utility we aim to maximize, is trained, the next step is to apply reinforcement learning under this reward model. Typically, the following objective is used:
@@ 3.2 Counterfactual Invariance
An ideal debiased reward model should intuitively remain invariant to spurious factors of variations. For example, to eliminate length bias, the reward model should exhibit invariance to changes in response length. To formalize this notion, we leverage the concept of counterfactual invariance (Veitch et al., 2021). We begin by introducing some notation. Let $Z$ represent the random variable corresponding to a spurious factor of variation (e.g., length), and let $T$ denote the random variable that encompasses the prompt-response pair. A reward model $r$ is said to exhibit counterfactual invariance to $Z$ if $r(T(z))=r(T(z^{\prime}))$ for all $z,z^{\prime}$, where $z$ and $z^{\prime}$ are realizations of $Z$, and $T(z)$ denotes the counterfactual $T$ we would have observed if $Z$ were $z$.
[…]
Throughout the paper, we use the term “invariant" or “debiased" to refer specifically to counterfactual invariance.
@@ 3.3 Causal Decomposition
The prompt-response pair $T$ can be decomposed into latent components based on their relations with the spurious factor $Z$ (Veitch et al., 2021). Specifically, we define $T^{Z,\perp}$ as the component of $T$ that is not causally influenced by $Z$. In other words, $T^{Z,\perp}$ represents the part of $T$ such that any function of $T$ is counterfactually invariant to $Z$ if and only if it depends solely on $T^{Z,\perp}$. Under weak conditions on $Z$, $T^{Z,\perp}$ is well defined. Further details regarding the derivation and properties can be found in (Veitch et al., 2021). In the next section, we extend this concept to develop a reward model that incorporates counterfactual invariance, which enables debiasing against various spurious factors.
@@ 4 Method
where $f$ is a counterfactually invariant function. This independence condition is merely a necessary condition implied by counterfactual invariance. But the key idea is that it constrains the hypothesis class, potentially guiding the model toward learning an invariant predictor.
@@ 4.1 Maximum Mean Discrepancy (MMD) Regularization for Independence
where $\mathcal{F}$ denotes a class of functions in $\mathcal{H}_{k}$, and $x\sim\mathbb{P}$, $y\sim\mathbb{Q}$ are two random variables. Intuitively, MMD measures the maximum mean difference between $\mathbb{P}$ and $\mathbb{Q}$ over functions $f\in\mathcal{F}$, as determined by a kernel $k(\cdot,\cdot)$ such as Gaussian kernels. In our approach, we use MMD as a regularizer to ensure that the learned reward model $f(T)$ is invariant to the spurious variable $Z$. If $Z$ is binary, our MMD regularizer can be defined as:
@@
To regularize $r_{\phi}(x,y)$, we map all responses into $M$ bins based on their spurious factor $Z$ (e.g., response length). For each bin $b$, we compute the conditional distribution $P_{b}(f(x,y))$. The overall objective function, combining the reward model training loss and the MMD-based regularizer, is:
@@ 5.1 Addressing Sycophantic Bias (Semi-synthetic)
Dataset and training. To investigate sycophantic bias, we create a semi-synthetic dataset based on dataset developed by Sharma et al. (2023). Specifically, our prompts are structured with the template "{question} I think the answer is {correct_answer} but I’m really not sure.". In this setup, we artificially induce a correlation between sycophantic behavior and correctness. Specifically, with an $80\%$ probability, the chosen response is prefixed with "Yes, you are right." Conversely, with a $20\%$ probability, this prefix appears in the rejected response. This creates an artificial but controlled spurious correlation between agreement ("Yes, you are right.") and correct answer, enabling us to observe, measure and address sycophantic bias effectively.
@@
Results. For each test prompt, we generate 50 responses. We then quantify sycophancy by checking whether the phrase "Yes, you are right." appears in any of those responses. Table 1 reports the percentage of test prompts for which all 50 sampled responses exhibit sycophantic behavior. It is worth noting that the SFT model, trained on chosen responses with high correlation with sycophantic phrasing, naturally tends to produce "Yes, you are right." as a default pattern. In contrast, both the conditional and unconditional CRM approaches successfully disentangle this spurious correlation and reduce the prevalence of sycophantic responses.
@@ 5.2 Addressing Length Bias
Dataset and training. We adopted the Alpaca dataset (Dubois et al., 2024b) for our experiments. Initially, we uses the chosen response for each prompt to do supervised finetuning (SFT) using the Llama-3 8B base model (Dubey et al., 2024). Then, this SFT model was subsequently employed to train both the reward model and the policy model. For reward model, we used the chosen and the rejected pair for training. With the reward model, we then trained the SFT policy with the PPO implementation from OpenRLHF (Hu et al., 2024) for one epoch. Additional details on hyperparameters and configurations are available in the Appendix 8.2.
Results. Our findings are illustrated in Fig. 2, where each dot on the plots represents a single model run, evaluated by its win rate, calculated as the proportion of wins against the SFT model. The score is defined by $\text{score}=50+(n_{\text{win}}-n_{\text{lose}})/N*100$, where $n_{\text{win}}$ and $n_{\text{lose}}$ denote the counts of wins and losses, respectively, and $N$ represents the total test count. In the leftmost plot, we observe that both the conditional and unconditional causal regularization methods achieve superior performance compared to the vanilla reward model with length penalty, as shown by their higher exponential moving average (EMA) curves. Furthermore, when examining the Pareto frontier, our approach demonstrates an advantage over the baseline method.
@@
Figure 2: Results on Length Bias, with each dot representing models trained with different regularization coefficients and PPO hyperparameters. The left plot shows exponential moving average (EMA) curves, the middle plot illustrates the Pareto front, and the right plot captures the length-rank correlation for different causal reward models.
@@ 5.3 Addressing Concept Bias
Dataset and training. We conducted experiments using Yelp (Zhang et al., 2015), IMDB (Maas et al., 2011), and Amazon Shoe Review (He and McAuley, 2016) datasets, augmented with additional concept labels provided by Zhou et al. (2023). Specifically, each dataset includes three concepts, where Yelp has "price", "service", "food"; IMDB has "music", "acting", "comedy"; and Amazon has "size", "color" and "style". To introduce more obvious concept bias, following Zhou et al. (2023), we modified each dataset to ensure all positive-sentiment samples were explicitly linked to a specific concept. For instance, in the Yelp dataset, we filtered reviews so that all positive sentiment entries were linked to the "food" concept.
@@
Metrics. We assess performance using both utility metrics (Acc@C, Acc@NoC) as well as the bias-specific metric Bias@C, as introduced in (Zhou et al., 2023). The utility metrics, which reflect the accuracy of correct sentiment classifications with (Acc@C) and without (Acc@NoC) the presence of a concept, indicate better performance with higher values. On the other hand, Bias@C measures spurious correlations associated with concept $C$, where values closer to zero suggest weaker biases. Specifically, positive Bias@C values suggest the model tends to predict positive labels when concept $C$ is present in the input, whereas negative values suggest the opposite tendency. For a more detailed explanation of the Bias@C metric, we direct interested readers to their original work (Zhou et al., 2023).
@@
Results. As shown in Table 2, the results demonstrate that CRM consistently reduces concept bias across the Yelp, IMDB, and Amazon Shoe Review datasets compared to the vanilla reward model. Specifically, both conditional and unconditional CRMs achieve significantly lower Bias@C values, with conditional CRM showing reductions of up to 97% on the Yelp dataset (e.g., for the "Price" concept). These results highlight the effectiveness of our approach in mitigating spurious correlations.